ninajaen:

Nick had 14 jelly beans in a bag. His mother put a handful of jelly beans in the bag. When Nick counts his jelly beans, he discovers he now has 28 of them. Which of the following equations will help Nick solve for the number of jelly beans, j, that his mom put in the bag? (3 points) a 14j = 28 b 14 + 28 = j c 14 + j = 28 d j − 14 = 28

Vocaloid:

he starts with 14 and his mother adds some amount j to get 28 thinking about that in terms of an equation, which one do you think best describes this?

Piffi:

to do this, read the first sentence. nick starts off having 14 jellybeans. write 14 down. then his mother ADDS in some jellybeans, but we don't know how much "some" is, so put down "+j." He has 28 jellybeans in TOTAL after his mother put in the "some" amount of jellybeans. Put down "=28."

Piffi:

so what you should have right now is: 14 + j = 28.

yourdootcome:

if you still dont understand then 14 + j = 28 is also equal to j = 28 - 14